OpenStudy (anonymous):

PLEASE HELP!! Adam wishes to travel from Paradise, to Cove located 500 km N60degreesW. The groundspeed of the plane is 380 km/h. The wind is from S80degreesE at 85 km/h. a) In what direction should Adam steer?b) What will the plane's airspeed be?c) How long will the trip take?

OpenStudy (anonymous):

I am so confused, heeeelp D:

OpenStudy (anonymous):

Hi, so all of these directions are vectors. Have you drawn a scale drawing?

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't know if mine is correct but here's what I've done: g=p+w p=g-w g = 380 km/h, [N60degreesW] p = ? w = 85 km/h [N80degreesW] 1. p = g-w = 380 km/h [N60degreesW] - 85 km/h [N80degreesW] = 380 km/h [N60degreesW] + 85 km/h [S80degreesE]  \[b ^{2} = a ^{2} + c ^{2} - 2accosB\] \[b ^{2} = 85^{2} + 380^{2} - 2(85)(380)\cos40\] \[= 151625-64600\cos40\] \[b = 258.2 km/h\]